OpenStudy (anonymous):

jackie has grades of 90 and 78 on her first two algebra tests. if she wants at least an average of 78 what possible score can she make on her thrid test? a) greater than 66 b)less than 66 c)66 or greater d)66 or less

OpenStudy (anonymous):

@aurorablue01

OpenStudy (anonymous):

i thinks its A

OpenStudy (anonymous):

Let x represent the unknown score. (90 + 78 + x)/3 = 78. Solve that--(168+x)/3 = 78 then, 168 + x = 234 Then subtract 168 from each side. x = 66 The answer is C. A is close but not quite, notice that 66 is a good enough score to get her an average of 78.